What is the probability of getting a license plate that has a repeated letter or digit if you live in a state that has three numerals followed by three letters followed by two numerals?
Answer by stanbon(57387)

What is the probability of getting a license plate that has a repeated letter or digit if you live in a state that has three numerals followed by three letters followed by two numerals?
-------------------------
P(at least one repeat) = 1 - P(no repetitions)
---
= 1 - (10*9*8*26*25*24)/(10^3*26^3)
---
= 1 - 0.6391
----
= 0.3609
==================
Cheers,
Stan H.
==================